A company having a net working capital of Rs 2.8 Lakhs as on 30-6-2018 indicates thefollowing financial ratios and performances figures:Current ratio 2.4Liquidity ratio 1.6Inventory turnover (on cost of sales) 8Gross profit on sales 20%Credit allowed (months) 1.5The company's fixed assets is equivalent to 90% of its net-worth (share capital plus reserve)while reserves amounted to 40% of share capital.Prepare the balance sheet of the company as on 30-6-2018 showing step by step calculations.

1
Expert's answer
2021-03-16T08:57:33-0400


find the missing data

"Current ratio=\frac{current assets}{current liabilities}"

"2.4=\frac{2.8}{current liabilities}"


current liabilities=1.2


"Liquidity ratio=\frac{Cash + Accounts Receivables + Marketable Securities}{Current Liabilities}"


"1.6=\frac{Cash + Accounts Receivables + Marketable Securities}{1.2}"

Cash + Accounts Receivables + Marketable Securities=1.92


Inventory=current assets-Cash + Accounts Receivables + Marketable Securities=2.8-1.92=0.88

"Inventory turnover =\frac{cost of sales}{Inventory}"

"8 =\frac{cost of sales}{0.88}"

cost of sales=7.04


let the share capital be 4

"net-worth (share capital plus reserve)=4+0.4\times4=5.6"

"fixed assets=0.9\times net-worth (share capital plus reserve)=0.9\times5.6=5.04" net-worth (share capital plus reserve)


the balance sheet of the company as on 30-6-2018:

fixed assets - 5.04

current assets - 2.8

total assets - 7.84


net-worth (share capital plus reserve) - 5.6

retained earnings - 1.04

current liabilities - 1.2

total  liabilities- 7.84